hartnn (hartnn):

Refreshing Differential Equations: How do i solve \( \Large e^{-x} sin y-e^{-x} cos y ~ y'-x+yy'=0 \)

hartnn (hartnn):

\(\Large e^{-x} sin y-e^{-x} cos y ~ y'-x+yy'=0\)

OpenStudy (usukidoll):

huh???

OpenStudy (usukidoll):

is this Bernoullis? cuz I haven't learned it yet.

hartnn (hartnn):

there is a sin y in there...so i don't think its bernoullis

OpenStudy (usukidoll):

\[-e^{-x}cosy y' +yy' +e^{-x}siny-x = 0\]

OpenStudy (usukidoll):

\[y'(-e^{-x}cosy+y)+e^{-x}siny-x=0\]

OpenStudy (usukidoll):

integrating factor???

OpenStudy (usukidoll):

oh wait no...

OpenStudy (usukidoll):

separable?!

OpenStudy (usukidoll):

exact? oh gawd not substitution D:

hartnn (hartnn):

not separable. can't think of any substitution as of now...

OpenStudy (usukidoll):

hmmm for a first order ode there are 5 methods, integrating factor, separable, substitution, exact, and ...... ugh I need to search for one more. I am getting a brain freeze here.

hartnn (hartnn):

i will check whether its exact or not..

OpenStudy (usukidoll):

homogeneous x+!_@

OpenStudy (usukidoll):

wait I got a supplement book... lists everything

OpenStudy (usukidoll):

nevermind.. it had what I listed

hartnn (hartnn):

woah, yes its exact

OpenStudy (anonymous):

Looks exact.

OpenStudy (usukidoll):

exact equation

hartnn (hartnn):

so, now finding Integrating factor

OpenStudy (usukidoll):

because integrating requires the y' +p(x)y=q(x) and there wasn't any y so forget that

OpenStudy (usukidoll):

don't you have to take the antiderivative in terms of x and take the derivative in terms of y

OpenStudy (usukidoll):

then whatever it was + h'(y) = N and something must cancel... and then take the antiderivative.

hartnn (hartnn):

wouldn't i just do \(\int M dx \\\int N dy\) ? and then combine those ?

OpenStudy (usukidoll):

what?

OpenStudy (anonymous):

\[ u = \exp \left( \int \frac{{M_y-N_x} }{N} dx \right) \]

OpenStudy (usukidoll):

oh wait ... are we trying to find the integrating factor first and then multiply it all the way through so when we take the partial derivatives it will be exact and then we do whatever I wrote earlier.

OpenStudy (anonymous):

That is one option.

OpenStudy (anonymous):

\[M(x,y)dx + N(x,y)dy = 0\]

hartnn (hartnn):

Mdx + Ndy = 0 thats the type of DE, i have now. and it is exact so don't i just integrate M and N, and then combine ? why do i even need to find Integrating factor! its already exact lol

OpenStudy (usukidoll):

Take yo M right... TAKE ANTIDERIVATIVE IN TERMS OF X AND TAKE THE DERIVATIVE IN TERMS OF Y! WEEEEEEEEEEEEEEEEEEEE

OpenStudy (anonymous):

Okay, if it is exact, you still need to find \(f\)

OpenStudy (anonymous):

What did you get for \(f\)?

hartnn (hartnn):

where is f ? :O are you asking for final answer ?

OpenStudy (kainui):

I feel like these methods sort of hide the fact that you're really just reverse-engineering the derivative.\[\Large \frac{d \Psi}{dx}=\Psi_x \frac{dx}{dx}+\Psi_y \frac{dy}{dx}\] It makes it easier to see what you have to integrate with respect to what.

OpenStudy (usukidoll):

:/

OpenStudy (anonymous):

\[ \Large (y-e^{-x} \cos y )~ dy + ( e^{-x} \sin y -x)~dx=0 \]

OpenStudy (anonymous):

@hartnn When I say 'looks exact' I mean 'looks like making it exact could work'. Integrating factor makes it exact.

OpenStudy (usukidoll):

 yeah I got this without integrating factor... doesn't look exact